# PepTune: De Novo Generation of Therapeutic Peptides with Multi-Objective-Guided Discrete Diffusion

### v1

@@ 2.4 Therapeutic Property Prediction for Peptide SMILES
While several classifiers exist for predicting properties of small-molecule SMILES sequences and amino-acid representations of peptides, there exists a gap in high-quality property models trained specifically on peptide SMILES data. To fill this gap, we train regression models for target-binding affinity and permeability and binary classification models for solubility, hemolysis, and non-fouling specifically on peptide SMILES data (Table 4).
@@
To guide the generation of peptides with high binding affinity to a given protein target, we trained a Transformer-based model with cross multi-head attention layers that learn the joint latent space of ESM-2-650M [44] embeddings of the protein amino acid sequence and PeptideCLM [45] embeddings of the peptide SMILES sequence (Figure 4. See full architectural details in Appendix C.2). Given a peptide SMILES sequence and a protein amino acid sequence, the model was trained on non-canonical and canonical peptide SMILES $K_{d}/K_{i}/IC50$ binding affinity to predict a score that indicates weak binding (< 6.0), medium binding (6.0-7.5), and high binding ($\geq$ 7.5). Our regression model achieved a strong Spearman correlation coefficient of 0.869 on the training data and 0.633 on the held-out validation data.
@@
We also trained a cell membrane permeability regression model by training an XGBoost [46] boosted tree model on PeptideCLM [47] embeddings which outputs a PAMPA score (logP) given a peptide SMILES sequence, where values $\geq-6.0$ indicate strong permeability and values $<6.0$ indicate weak permeability. We trained our model on 34,853 experimentally validated peptide SMILES (See Methods 4.1) and achieved a strong Spearman correlation coefficient of 0.998 on the training dataset and 0.943 on the test dataset (Figure 5, Table 5).
@@ 2.5 Case Studies for Multi-Objective Generation of Peptide Binders
With our trained property classifiers, we first conduct experiments for five diverse, therapeutically relevant protein targets to evaluate our multi-objective MCTS guidance strategy. To demonstrate generalizability, we include targets with known peptide binders such as GLP-1R and TfR and proteins with no known binders, including GFAP, GLAST, and AMHR2. These targets include both receptor proteins involved with active transport pathways as well as intracellular targets where cell membrane permeability is crucial to achieving therapeutic effects. For each target, we condition the generation on the binding affinity score given the target protein sequence along with solubility, hemolysis, non-fouling, and cell membrane permeability for intracellular targets. For external testing and validation, we use Autodock Vina [48] to compute in silico binding affinities of our generated binders (See Methods 4.5).
Targeting Receptors on the Blood-Brain Barrier. The Transferrin receptor (TfR) is a receptor protein abundant on the selectively permeable blood-brain barrier (BBB) that is responsible for transporting iron-binding transferrin proteins into the brain parenchyma [49]. Given its selective expression on brain endothelial cells and glioma cells and its ability to recycle back to the luminal surface after facilitating the internalization of cargo through the BBB [50], TfR has been extensively studied as a target for the intravenous delivery of various therapeutics and therapeutic nanocarriers through the BBB [51, 52].
@@
Figure 6: PepTune-generated peptide binders to TfR. (A) Density plot depicting the frequency of predicted binding affinity scores from our trained regression model for the sequences in the data used to train the regression model, the generated peptides from our unconditional PepMDLM model, and our PepTune model conditioned on TfR binding affinity, solubility, hemolysis, and non-fouling. (B) Plots depicting the mean scores for each property over the number of iterations or traversals of the MCTS algorithm for 128 iterations and a maximum token length of 200. The shaded region represents the standard deviation. (C) Two-dimensional visualization of generated binders with token length 100, their corresponding docking scores ($\downarrow$) computed using Vina docking, and predicted scores ($\uparrow$) from the trained classifiers. (D) Visualizations of generated binders with token length 200, their docking scores, and predicted classifier scores.
To further confirm binding affinity to TfR, we compared our peptides to the well-established 7-amino acid peptide T7 (sequence: HAIYPRH) that selectively binds to an alternative site as compared to endogenous Tf on TfR [53]. T7 has been extensively explored for targeted delivery of nanoparticles to the brain [54, 55, 56, 57, 58, 59, 60], and has demonstrated 7.89-fold enhanced brain penetration in in vivo mice models [61]. After docking T7 with TfR, we obtained a docking score of -8.4 kcal/mol. Notably, our peptides optimized on all four therapeutic properties including TfR binding affinity show competitive docking scores to T7 (Figure 7A, C, E), suggesting that PepTune is capable of generating promising candidates for future in vivo delivery across the BBB. Furthermore, after annotating polar contacts within 3.5 Å  we determine that both of the generated peptides with the best binding affinity scores have shared residue contacts when binding with TfR as T7 (Figure 7B, D, F), indicating that our generated peptides have superior binding properties to T7, enabling it to bind more strongly with the target despite a shared binding site. Furthermore, our generated binders have diverse structural features, such as cycles in binder 1 and side-chain modifications in binder 2. Since T7 is known to bind to an alternative site than endogenous Tf [53], we show that PepTune can generate viable candidates for non-competitive binding to TfR for enhanced targeting.
@@
Shorter agonists or antagonists to GLP-1R would serve several benefits to the treatment of insulin-related disorders, including reduced cost and complexity of synthesis, lower immunogenicity, and faster tissue penetration. Therefore, we sought to generate shorter-chain peptides that are capable of binding to GLP-1R with comparable affinity to the existing agonists. We first generated a pool of peptide binders conditioned on binding affinity with the GLP-1R sequence, solubility, hemolysis, and non-fouling. After selecting the peptides with the highest affinity scores from the Pareto-non-dominant set, we performed docking and determined affinities of -7.4 kcal/mol and -7.0 kcal/mol for the two best candidates. We note that our peptides have strong affinity to GLP-1R and while also interacting at motifs that strongly overlap with the binding sites of semaglutide and liraglutide derived from its natural hormone ligand, GLP-1 (Figure 8). These results suggest that our PepTune-derived peptides can serve as potent agonists or antagonists of GLP-1R signaling.
@@
Targeting Intracellular Proteins. Glial fibrillary acidic protein (GFAP) is an intracellular protein differentially expressed in astrocytes, a family of glial cells in the brain [64]. Dysregulation of GFAP expression has been found to cause Rosenthal fibers, astrocytic cytoplasmic inclusions that are responsible for Alexander disease, a fatal neurodegenerative disease affecting infants [65, 66]. Discovering potent binders that inhibit or degrade GFAP proteins can have significant therapeutic implications. However, no established peptide binders exist to GFAP, which motivates their de novo design. In addition to achieving high binding affinity with GFAP, we posit that an optimal peptide binder must also cross the astrocyte cell membrane into the cytosol to access GFAP. Therefore, we condition the generation of GFAP binders on five properties: binding affinity to GFAP, solubility, hemolysis, non-fouling, and cell membrane permeability using our permeability regression model, demonstrating optimization across all of these properties (Figure 9). To confirm GFAP engagement, our docking results demonstrate strong affinities of our designed peptides, motivating downstream experimental validation in astrocyte cultures (Figure 9B and D).
@@
Figure 9: PepTune-generated peptide binders to intracellular protein GFAP. (A, C) Two-dimensional structures of GFAP binder 1 and 2 with predicted property scores, including cell membrane permeability. (B, D) GFAP binders 1 and 2 docked to GFAP with score of -8.5 kcal/mol and -7.1 kcal/mol respectively. (E) Full GFAP protein structure with docked binders 1 and 2. (F) The distribution of PAMPA membrane permeability scores from 34,853 experimentally-validated peptides compared to 100 peptides generated using our unconditional PepMDLM model, and 100 peptides generated with PepTune conditioned on both cell membrane permeability and affinity to GFAP. The permeability curve shifted towards higher permeability with mean of -6.295. (G) Simultaneously, the distribution of predicted binding affinity scores to GFAP for the PepTune-generated peptides is shifted to higher scores with a mean of 8.053 compared to a set of experimentally-tested peptides and unconditional PepMDLM-generated peptides.
@@
Targeting Extracellular Proteins Without Existing Binders. To test the ability of our model to generate binders to challenging extracellular targets without existing binders, we evaluate PepTune-generated peptides for NCAM1 and AMHR2, two therapeutically-relevant receptor proteins. Neural cell adhesion molecule 1 (NCAM1), is a transmembrane protein expressed on the surface of neurons and glial cells [67]. Beyond its roles in neuronal migration and synaptogenesis, NCAM1 is also crucial for memory formation, highlighting its significance in brain development [68]. As NCAM1 is an extracellular protein, we generated a library of peptides with PepTune optimized NCAM1 binding affinity, solubility, hemolysis, and non-fouling (Figure 10F, G). All properties exhibited an upward trend across optimization iterations.
@@
2.6 PepTune Generates Multi-Target Specific Peptides
@@
PepTune is uniquely positioned to tackle the multi-target optimization task since it can explore several different unmasking pathways while maintaining a set of Pareto-optimal peptide sequences with non-dominated binding affinity scores with each of the protein targets. Our strategy enables conditioning on multiple target proteins to design binders with high affinity to both targets without sacrificing the discovery of peptides that bind strongly to only one of the targets since the model will return all non-dominant peptides.
@@
Targeting TfR and GLAST for Drug Delivery to Astrocytes. To evaluate PepTune’s capabilities in multi-target guidance, we use PepTune to generate bi-specific peptide binders to TfR and glutamate-aspartate transporter (GLAST) protein abundant on the surface of astrocytes, a type of glial cell in the brain. Successfully generating these peptides can facilitate BBB-crossing via TfR binding and uptake in astrocytes via GLAST binding for intravenous delivery of therapeutics for a multitude of neurological disorders where astrocytes are involved, including Alexander disease [77], Alzheimer’s disease [78], Parkinson’s disease [79], Huntington’s disease [80], multiple sclerosis [81], and several psychiatric disorders [82].
We generated a pool of 100 peptide binders conditioned on five properties: predicted binding affinity to TfR, predicted binding affinity to GLAST, solubility, hemolysis, and non-fouling. Notably, we observed an increase across all properties over iterations, with the final solubility, hemolysis, and non-fouling scores surpassing the binders conditioned only on TfR binding affinity (Figure 12). This suggests that multi-target guidance does not result in significant trade-offs in property scores.
To confirm that our generated binders indeed bind to both TfR and GLAST, we selected seven binders and conducted docking against TfR and GLAST separately for each binder. Incredibly, the docking scores across all seven binders were less than or equal to -7.5 kcal/mol for both targets, with the best-scoring binder simultaneously achieving a score of -10.5 kcal/mol for TfR and a score of -9.2 kcal/mol for GLAST (Table 6). In addition, the top-performing binders have diverse secondary structures (Figure 13) and have positive solubility, and hemolysis probabilities (Table 6). The binding positions and polar interactions vary greatly across the top-performing candidates, enabling the selection of binders that fit specific location constraints. This indicates that PepTune can discover a wide subspace of optimal peptides with strong binding affinity to both TfR and GLAST that is not dependent on a specific binding motif. Our next steps consist of validating the dual-binding affinity of our top candidates in an in vitro BBB-transwell model and observing whether both BBB-crossing and uptake into basolateral astrocytes are enhanced for dual-target compared to single-target conditioned and control peptides.
@@
Dual-Targeting of GFAP and an E3 Ubiquitin Ligase for Target Protein Degradation. As another dual-target case study, we used PepTune to generate peptides with high binding affinity to GFAP protein and an E3 ubiquitin ligase protein RBX1, a protein in the Skp1/Cullin-1/F-box (SCF) E3 ubiquitin ligase complex that recruits E2 to catalyze ubiquitination [83]. A peptide generated for this task would be capable of binding to GFAP proteins overexpressed in Alexander disease and mediating their proteasomal degradation, which could alleviate the production of disease-causing Rosenthal fibers in astrocytes [84].
@@
After conditioning PepTune generation on binding affinity to GFAP, binding affinity to RBX1, solubility, hemolysis, and non-fouling (Table 7), we selected three non-dominated binders with predicted affinities greater than 7.0 for docking experiments. For these PepTune-generated peptides, we indeed observed strong binding affinities for both GFAP and RBX1 post-docking, indicating their unique potential for multi-target interaction (Figure 14). As a note, GFAP is an intermediate filament protein [85], and thus forms a unique rod-like structure with a head domain and a tail domain. The docking positions of all three candidates were along the rod domain, binding in the gap between adjacent rods in the filament. Contrarily, docked candidates to RBX1 consistently bound close to its interaction site of Cullin, rather than at the Skp2 F-box adaptor site (Figure 14), indicating that further motif conditioning, as done with recent peptide design language models [86], would benefit PepTune’s clinical potential.

### v2

@@ 2.4 Therapeutic Property Prediction for Peptide SMILES
While several classifiers exist for predicting properties of small-molecule SMILES sequences and amino-acid representations of peptides, there exists a gap in high-quality property models trained specifically on peptide SMILES data. To fill this gap, we train regression models for target-binding affinity and cell membrane permeability and binary classification models for solubility, hemolysis, and non-fouling specifically on peptide SMILES data (Table 4).
@@
To guide the generation of peptides with high binding affinity to a given protein target, we trained a Transformer-based model with cross multi-head attention layers that learn the joint latent space of ESM-2-650M [44] embeddings of the protein amino acid sequence and PeptideCLM [45] embeddings of the peptide SMILES sequence (Figure 4. See full architectural details in Appendix C.3). Given a peptide SMILES sequence and a protein amino acid sequence, the model was trained on natural and non-natural peptide SMILES $K_{d}/K_{i}/IC50$ binding affinity to predict a score that indicates weak binding (< 6.0), medium binding (6.0-7.5), and high binding ($\geq$ 7.5). Our regression model achieved a strong Spearman correlation coefficient of 0.869 on the training data and 0.633 on the held-out validation data.
@@
For cell membrane permeability, we trained an XGBoost [46] boosted tree regression model on PeptideCLM [47] embeddings which returns the predicted PAMPA lipophilicity score ($\log P$) given a peptide SMILES sequence, where values $\geq-6.0$ indicate strong permeability and values $<6.0$ indicate weak permeability. We trained our model on 34,853 experimentally validated peptide SMILES (See Methods 4.1) and achieved a strong Spearman correlation coefficient of 0.998 on the training dataset and 0.943 on the test dataset (Figure 5, Table 5).
@@ 2.5 Case Studies for Multi-Objective Generation of Peptide Binders
With our trained property classifiers, we first conduct experiments for five diverse, therapeutically relevant protein targets to evaluate our multi-objective MCTS guidance strategy. To demonstrate generalizability, we include targets with known peptide binders such as GLP-1R and TfR, and proteins with no known binders, including GFAP, NCAM1, and AMHR2. These targets include both receptor proteins involved with active transport pathways as well as intracellular targets where cell membrane permeability is crucial to achieving therapeutic effects. For each target, we condition the generation on the binding affinity score given the target protein sequence along with solubility, hemolysis, non-fouling, and cell membrane permeability for intracellular targets. For external testing and validation, we use Autodock Vina [48] to compute in silico binding affinities of our generated binders (See Methods 4.5).
Targeting Receptors on the Blood-Brain Barrier. The Transferrin receptor (TfR) is a receptor protein abundant on the selectively permeable blood-brain barrier (BBB) that is responsible for transporting iron-binding transferrin (Tf) proteins into the brain parenchyma [49]. Given its selective expression on brain endothelial cells and glioma cells and its ability to recycle back to the luminal surface after facilitating the internalization of cargo through the BBB [50], TfR has been extensively studied as a target for the intravenous delivery of various therapeutics and therapeutic nanocarriers through the BBB [51, 52].
@@
Figure 6: PepTune-generated peptide binders to TfR. (A) Density plot depicting the frequency of predicted binding affinity scores from our trained regression model for the sequences in the data used to train the regression model, the generated peptides from our unconditional PepMDLM model, and our PepTune model conditioned on TfR binding affinity, solubility, hemolysis, and non-fouling. (B) Plots depicting the mean scores for each property over the number of iterations or traversals of the MCTS algorithm for 128 iterations and a maximum token length of 200. The shaded region represents the standard deviation. (C) Two-dimensional visualization of generated binders with token length 100, their corresponding docking scores ($\downarrow$) computed using Vina docking, and predicted classifier scores ($\uparrow$) from the trained classifiers. (D) Visualizations of generated binders with token length 200, their docking scores, and predicted classifier scores.
To further confirm binding affinity to TfR, we compared our peptides to the well-established 7-amino acid peptide T7 (sequence: HAIYPRH) that selectively binds to an alternative site as compared to endogenous Tf on TfR [53]. T7 has been extensively explored for targeted delivery of nanoparticles to the brain [54, 55, 56, 57, 58, 59, 60], and has demonstrated 7.89-fold enhanced brain penetration in in vivo mice models [61]. After docking T7 with TfR, we obtained a docking score of -8.4 kcal/mol. Notably, our peptides optimized on all four therapeutic properties including TfR binding affinity show competitive docking scores to T7 (Figure 7A, C, E), suggesting that PepTune is capable of generating promising candidates for in vivo targeting and delivery across the BBB. Furthermore, after annotating polar contacts within 3.5 Å  we determine that both of the generated peptides with the best binding affinity scores have shared residue contacts when binding with TfR as T7 (Figure 7B, D, F), indicating that our generated peptides have similar binding properties to T7, enabling it to bind strongly to a shared binding site. Furthermore, our generated binders have diverse structural features, such as cycles in binder 1 and side-chain modifications in binder 2. Since T7 is known to bind to an alternative site than endogenous Tf [53], we show that PepTune can generate viable candidates for non-competitive binding to TfR for BBB-targeting applications.
@@
Shorter agonists or antagonists to GLP-1R would serve several benefits to the treatment of insulin-related disorders, including reduced cost and complexity of synthesis, lower immunogenicity, and faster tissue penetration. Therefore, we sought to generate shorter-chain peptides that are capable of binding to GLP-1R with comparable affinity to the existing agonists. We first generated a pool of peptide binders conditioned on binding affinity with the GLP-1R sequence, solubility, hemolysis, and non-fouling. After selecting the peptides with the highest predicted binding affinity scores from the Pareto non-dominated set, we performed docking and determined docking scores of -7.4 kcal/mol and -7.0 kcal/mol for the two best candidates. Our peptides show superior docking affinity to GLP-1R while interacting at overlapping binding motifs to semaglutide and liraglutide derived from the natural hormone ligand, GLP-1 (Figure 8). These results suggest that our PepTune-derived peptides can serve as potent agonists or antagonists of GLP-1R signaling.
@@
Targeting Intracellular Proteins. Glial fibrillary acidic protein (GFAP) is an intracellular protein differentially expressed in astrocytes, a family of glial cells in the brain [64]. Dysregulation of GFAP expression has been found to cause Rosenthal fibers, astrocytic cytoplasmic inclusions that are responsible for Alexander disease, a fatal neurodegenerative disease affecting infants [65, 66]. Discovering potent binders that inhibit or degrade GFAP proteins can have significant therapeutic implications. However, no established peptide binders exist to GFAP, which motivates their de novo design. In addition to achieving high binding affinity with GFAP, we posit that an optimal peptide binder must also cross the astrocyte cell membrane into the cytosol to access GFAP. Therefore, we condition the generation of GFAP binders on five properties: binding affinity to GFAP, solubility, hemolysis, non-fouling, and cell membrane permeability using our permeability regression model, demonstrating optimization across all of these properties (Figure 9). To confirm GFAP engagement, our docking peptides demonstrate strong affinities below -7 kcal/mol, motivating downstream experimental validation in astrocyte cultures (Figure 9B and D).
@@
Figure 9: PepTune-generated peptide binders to intracellular protein GFAP. (A, C) Two-dimensional structures of GFAP binder 1 and 2 with predicted property scores, including cell membrane permeability. (B, D) GFAP binders 1 and 2 docked to GFAP with score of -8.5 kcal/mol and -7.1 kcal/mol respectively. (E) Full GFAP protein structure with docked binders 1 and 2. (F) The distribution of PAMPA membrane permeability scores from 34,853 experimentally-validated peptides compared to 100 peptides generated using our unconditional PepMDLM model, and 100 peptides generated with PepTune conditioned on both cell membrane permeability and affinity to GFAP. The permeability curve shifted towards higher permeability with a mean of -6.295. (G) Simultaneously, the distribution of predicted binding affinity scores to GFAP for the PepTune-generated peptides is shifted to higher scores with a mean of 8.053 compared to a set of experimentally-tested peptides and unconditional PepMDLM-generated peptides.
@@
Targeting Extracellular Proteins Without Existing Binders. To test the ability of our model to generate binders to challenging extracellular targets without existing binders, we evaluate PepTune-generated peptides for NCAM1 and AMHR2, two therapeutically relevant receptor proteins. Neural cell adhesion molecule 1 (NCAM1), is a transmembrane protein expressed on the surface of neurons and glial cells [67]. Beyond its roles in neuronal migration and synaptogenesis, NCAM1 is also crucial for memory formation, highlighting its significance in brain development [68]. As NCAM1 is an extracellular protein, we generated a library of peptides with PepTune-optimized NCAM1 binding affinity, solubility, hemolysis, and non-fouling (Figure 10F, G). All properties exhibited an upward trend across optimization iterations.
@@
2.6 Case Studies for Dual-Target Binding Peptides
@@
PepTune is uniquely positioned to tackle the multi-target optimization task since it can explore several different unmasking pathways while maintaining a set of Pareto-optimal peptide sequences with non-dominated binding affinity scores with each of the protein targets. Our strategy enables conditioning on multiple target proteins to design binders with high affinity to both targets without sacrificing the discovery of peptides with superior binding affinity to only one of the targets since the model keeps track of all non-dominated peptides.
@@
Targeting TfR and GLAST for Drug Delivery to Astrocytes. To evaluate PepTune’s capabilities in multi-target guidance, we generate bi-specific peptide binders to TfR and glutamate-aspartate transporter (GLAST) protein abundant on the surface of astrocytes, a type of glial cell in the brain. Successfully generating these peptides can facilitate BBB-crossing via TfR binding and uptake in astrocytes via GLAST binding for intravenous delivery of therapeutics for a multitude of neurological disorders where astrocytes are involved, including Alexander disease [77], Alzheimer’s disease [78], Parkinson’s disease [79], Huntington’s disease [80], multiple sclerosis [81], and several psychiatric disorders [82].
We generated a pool of 100 peptide binders conditioned on five properties: predicted binding affinity to TfR, predicted binding affinity to GLAST, solubility, hemolysis, and non-fouling. Notably, we observed an increase across all properties over iterations, with the final solubility, hemolysis, and non-fouling scores surpassing the binders conditioned only on TfR binding affinity (Figure 12). This suggests that our multi-target guidance strategy does not result in significant trade-offs in property scores.
To confirm that our generated binders indeed bind to both TfR and GLAST, we selected seven binders and conducted docking against TfR and GLAST separately for each binder. Incredibly, the docking scores across all seven binders were less than or equal to -7.5 kcal/mol for both targets, with the best-scoring binder simultaneously achieving a score of -10.5 kcal/mol for TfR and a score of -9.2 kcal/mol for GLAST (Table 6). In addition, the top-performing binders have diverse secondary structures (Figure 13) and have positive solubility, and hemolysis probabilities (Table 6). The binding positions and polar interactions vary greatly across the top-performing candidates, enabling the selection of binders that fit a range if binding site constraints. This indicates that PepTune can discover a wide subspace of optimal peptides with strong binding affinity to both TfR and GLAST that is not dependent on a specific binding motif. Our next steps consist of validating the dual-binding affinity of our top candidates in an in vitro BBB-transwell model and observing whether both BBB-crossing and uptake into basolateral astrocytes are enhanced for dual-target compared to single-target conditioned and control peptides.
@@
Dual-Targeting of GFAP and an E3 Ubiquitin Ligase for Target Protein Degradation. As another dual-target case study, we used PepTune to generate peptides with high binding affinity to GFAP protein and an E3 ubiquitin ligase protein RBX1, a protein in the Skp1/Cullin-1/F-box (SCF) E3 ubiquitin ligase complex that recruits E2 to catalyze ubiquitination and subsequent degradation [83]. A peptide generated for this task would be capable of binding to GFAP proteins overexpressed in Alexander disease and mediate their proteasomal degradation, which could alleviate the production of disease-causing Rosenthal fibers in astrocytes [84].
@@
After conditioning PepTune generation on binding affinity to GFAP, binding affinity to RBX1, solubility, hemolysis, and non-fouling (Table 7), we selected three non-dominated binders with predicted affinities greater than 7.0 for docking experiments. For these Pareto-optimal peptides, we indeed observed strong binding affinities for both GFAP and RBX1 post-docking, indicating their unique potential for multi-target interaction (Figure 14). GFAP is an intermediate filament protein [85] and thus forms a unique rod-like structure with a head domain and a tail domain. The docking positions of all three candidates were along the rod domain, binding in the gap between adjacent rods in the filament. Contrarily, docked candidates to RBX1 consistently bound close to its interaction site of Cullin, rather than at the Skp2 F-box adaptor site (Figure 14), indicating that further motif conditioning, as done with recent peptide design language models [86], would benefit PepTune’s clinical potential.